OpenStudy (anonymous):

1/(a-b) + 4/(b-a) - 8/(a+b) - (11a - 5b)/(b^2-a^2) does this equal (22a-10b)/(b^2-a^2)?

OpenStudy (anonymous):

I told you it was 0. Check all your signs.

OpenStudy (anonymous):

First, we make (a-b) = -(b-a): [1/(a-b) - 4/(a-b)] - 8/(a+b) - (11a - 5b)/(b^2-a^2) Subtracting the first two fractions: -3/(a-b) - 8/(a+b)) - (11a - 5b)/(b^2-a^2) Gathering the first two fractions into one fraction: (-3a - 3b - 8a + 8b)/(a-b)(a+b) - (11a - 5b)/(b^2-a^2) Equals: (-11a + 5b)/(a^2 - b^2) Doing the same as the first step: (11a - 5b)/(b^2 - a^2)

OpenStudy (anonymous):

Of course, note that: (b-a)(b+a) = (b^2 - a^2), right?